# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v2

@@ VII Mode overlaps
Thus far, we have assumed i.i.d. modes, with the factorization $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$ implying that left–right overlaps $\bm{\ell}_{a}^{T}\bm{r}_{a}$ are random and $\mathcal{O}(1/\sqrt{N})$. The same holds for $\bm{\ell}_{a}^{T}\bm{r}_{b}$ with $a\neq b$ due to the more fundamental assumption that $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ factorizes across $a$. However, computations in biological circuits require structured interactions between modes, suggesting stronger, non-random overlaps between left and right modes. To test this, we return to the Drosophila connectome and compute the overlap matrix between left and right singular vectors, $O_{ab}=\bm{u}_{a}^{T}\bm{v}_{b}$ (Fig. 6a). Contrary to the small, unstructured overlaps expected under the i.i.d. assumption, the connectome exhibits large, structured overlaps, particularly along the diagonal (Fig. 6b).
@@
Figure 6: Right-left mode overlaps and eigenvalue spectra of the fly-brain connectome. (a) Overlap matrix between right and left singular vectors of the connectome. First 1000 modes are shown. (b) First 3000 diagonal elements of the overlap matrix from (a). Horizontal lines indicate an approximate cutoff for $\mathcal{O}(1/\sqrt{N})$ random overlaps. (c) Eigenvalue spectra of various connectome-based matrices. Top: Actual hemibrain spectrum, exhibiting large outliers along the real axis. Middle: Eigenvalues from a realization of the random-mode model, using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum. Bottom: Random-mode model realization with right-left mode overlaps incorporated, where overlaps are given by the inner products of left and right singular vectors from the fly connectome SVD. Data from Scheffer et al. [17], $N=18028$ neurons.
@@
Motivated by the presence of these large diagonal overlaps, we extend the random-mode model to include correlations between corresponding left and right modes. Specifically, we take $P(\ell_{a},r_{a})$ to be a zero-mean bivariate Gaussian with marginal variance $1/N$ and covariance $\rho_{a}/N$, where $\rho_{a}$ is the correlation between the $a$-th mode pair. Both $D_{a}$ and $\rho_{a}$ are treated deterministically, and we assume that $M^{-1}\sum_{a}f(D_{a},\rho_{a})$ converges to a limiting value $\left\langle f(D,\rho)\right\rangle_{D,\rho}$ as $N\to\infty$. In this setting, $\bm{\ell}_{a}^{T}\bm{r}_{a}=\rho_{a}$ for large $N$, while $\bm{\ell}_{a}^{T}\bm{r}_{b}$ for $a\neq b$ remains random and $\mathcal{O}(1/\sqrt{N})$. This additional structure produces eigenvalue spectra resembling that of the connectome, including large real outliers (Fig. 6c) [59].
@@
This purely diagonal overlap structure, however, can only generate real eigenvalue outliers. The presence of complex outliers in the connectome, together with the full overlap matrix itself (Fig. 6a), implies the presence of richer correlations. Capturing such features would require relaxing the full factorization of $P(\{\ell_{a},r_{a}\})$ across $a$ while retaining $\mathcal{O}(M)$ parameter scaling—for example, via blockwise factorization or Markovian correlations across $a$. We leave such generalizations for future work. We also note that by setting $\rho_{a}=1$ for all $a$, the model becomes a Hopfield network storing an extensive set of Gaussian patterns. In such networks, each mode can be “condensed” ($\mathcal{O}(1)$ overlap with activity) or “uncondensed” ($\mathcal{O}(1/\sqrt{N})$ overlap with activity). Our analysis assumes no condensed modes, but sufficiently strong positive overlaps break this assumption. The Gaussian case, however, is unable to generate multiple multistable attractors as in a memory system [35]. Note that the form of the single-element density is relevant for the condensed patterns, but not for the uncondensed patterns. Allowing for condensed patterns within our analysis would require specifying this density. See [60] for a DMFT analysis of generalized Hopfield dynamics in which these ideas are implemented.
@@
In Appendix K, we compute the two- and four-point functions. As in previous cases, the dynamics reduce to single-site processes:
@@
To gain intuition for the self-coupling kernel $R_{\star}(\omega)$, we expand in powers of $\rho$. The coefficients encode connectivity motifs of increasing orders:
@@
We validate this theory using $\alpha=1$ and $D_{a}\propto\exp(-\beta_{D}a/N)$, as before. For overlaps, we use a sigmoidal spectrum for $\rho_{a}$ qualitatively inspired by Fig. 6b, with multiplicative prefactor $\gamma_{\rho}$ and other parameters chosen such that $\left\langle D\rho\right\rangle_{D,\rho}=0$ (Appendix K.3). Across values of $g_{\text{eff}}$ and $\text{PR}^{D}$, sufficiently strong overlaps, of either sign, reduce the dimension of activity (Fig. 7). The maximum dimension occurs for $\gamma_{\rho}<0$, where dynamics most strongly suppress activity aligned with dominant connectivity modes. This effect weakens as $g_{\text{eff}}$ increases.
Figure 7: Effect of right-left mode correlations on dimension of activity in the random-mode model. Dimension of activity $\text{PR}^{\phi}$ versus correlation parameter $\gamma_{\rho}$ in the $(D,\rho)$ joint distribution, for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$. Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions. All simulations use $N=5000$ neurons.
@@ VIII Discussion
Motifs vs. global structure. Neuronal connectivity is often described using statistics of motifs: local patterns like chains or cycles among small groups of neurons [62, 18, 63, 64]. These can be estimated from partial observations of a weight matrix which, until recently, set the limits of experimental measurement. Advances in whole-brain connectome reconstruction now allow access to global connectivity, including spectral features. Our parameterization of coupling matrices represents this global structure, with control over both the spectrum and mode overlaps. These spectral features determine how activity modes interact dynamically and thus may more directly dictate neural computations compared to motifs.
@@
That said, in some studies the primary goal is to capture motifs or graph-theoretic features (e.g., degree distributions) over which the random-mode model does not provide control (although some of these features have spectral signatures that could be incorporated). When such features are the focus, other connectivity models are better suited, for example maximum-entropy random graph models constrained by fixed structural features (e.g., the configuration model for matching degree distributions).
@@
Linking connectivity and activity datasets. The random-mode model, together with its dynamical theory, could aid in the interpretation of large-scale neural activity recordings in the context of anatomical connectivity, such as in Drosophila, where both the connectome and whole-brain activity data are available [65]. In the current work, we used this connectome to characterize the spectral structure of an actual large neural circuit, stopping short of considering the activity that it generates. Future work could progressively incorporate richer connectome constraints into the model to identify which features most strongly shape collective activity. One possible sequence would be to add the spectrum of component strengths, then diagonal mode overlaps, and finally off-diagonal overlaps, assessing at each step how much the correspondence between modeled and observed activity improves. This is similar to approaches in statistical physics that aim to recover higher-order correlations by constraining lower-order statistics [66, 67, 68].
@@
Learning. Learning alters collective dynamics by modifying connectivity—the core principle of machine learning. Our analysis focuses on structured but untrained networks that produce chaotic activity. Nevertheless, the extensive low-rank connectivity structures we study provide a starting point for understanding trained networks.
@@
Large networks trained on simple tasks often show intensive low-rank modifications to their initial couplings, with a finite number of outlier eigenvalues emerging from the bulk [23, 34]. Existing theoretical tools can already describe such systems by taking the trained weights, performing a low-rank decomposition, modeling mode overlaps statistically, and studying the resulting low-rank DMFT [36]. By contrast, empirical studies of large networks trained on real-world tasks suggest the emergence of extensive low-rank structure, with smooth spectral profiles with many significant components [24]. This structure resembles that modeled by the random-mode model. Whether mode overlaps extracted from such networks can be incorporated into the random-mode model framework, and thereby yield a working theory for the extensive-rank case, is an open question.
@@
A more ambitious direction is to develop a DMFT for networks whose weights are learned from data, rather than first training networks and then retrofitting a statistical model. Such a theory would describe not only how data shapes weights, as addressed by classical works [69], but also how data interacts with recurrent network dynamics. The path-integral formalism is well-suited to this, as it allows derivation of DMFT equations for coupling matrices drawn from a Gibbs ensemble with an energy function given by a task loss. While this approach has been applied to low-dimensional tasks, leading to intensive-rank weights [70], applying it to high-dimensional tasks could yield a theory of how such tasks generate extensive-rank weights that in turn produce extensive-dimensional activity whose single-neuron and collective properties the theory would describe.
@@
Another route to extensive-rank structure is via networks that perform many tasks, which could be modeled, for example, by representing mode overlaps through blockwise factorization of $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ across $a$. This “multi-task” scenario is increasingly relevant in both neuroscience and machine learning [71, 72].
@@
Even in standard RNN training, initializing couplings with low-rank structure could improve procedures such as FORCE learning, which rely on suppressing chaos [12, 13]. Lower-dimensional attractors may be easier to control, potentially leading to faster convergence.
@@
Modeling functional consequences of single-neuron heterogeneity. A central question in neuroscience is how neuronal diversity shapes collective dynamics and computation [73]. Appendix J presents a general framework that assigns each neuron a parameter vector $\bm{\theta}$ and enables study of how different forms of heterogeneity affect properties such as dimensionality. Sec. VI focuses on gain heterogeneity to model firing-rate diversity, but this framework could also incorporate more intricate cell-specific features, such as electrophysiological characteristics [74], linking single-neuron heterogeneity to functional consequences for population activity.
@@
Transient dynamics and excitation–inhibition. Our analysis has focused on stationary states with translation-invariant temporal correlations. Many neural computations, however, rely on transient dynamics that break stationarity and depend on absolute time. Extending the DMFT accordingly is numerically tractable for two-point functions but would become unwieldily for four-point functions, $\Psi^{a}_{\star}(t_{1},t_{2},t_{3},t_{4})$. Representing and manipulating such a large object may require approximations or low-rank decomposition.
@@
One notable transient phenomenon is transient amplification, linked to non-normal coupling matrices that naturally arise from excitatory–inhibitory segregation (Dale’s law) [75, 76]. Although the random-mode model does not enforce Dale’s law, the resulting matrices are nevertheless non-normal and may exhibit transient amplification; such effects may be amplified through structured off-diagonal overlaps between left and right modes. Non-normality also decouples the highest-variance dimensions from the slowest modes. Relatedly, extending the random-mode model to respect Dale’s law and incorporate a non-saturating $\phi(x)$—achieving stability via excitation–inhibition balance rather than saturation [77, 78]—could enable more direct comparisons to both connectivity and activity data.
@@
Random basis property. The random-mode model assumes that the $2M$ mode components associated with each neuron are sampled i.i.d. for each neuron (Eq. 3). This leads to neuronal permutation symmetry of the distribution over $\bm{J}$. Furthermore, when the marginal distributions over left and right mode components are isotropic Gaussians, as we have assumed, the embedding into neuronal space is, by definition, random and Gaussian. This “random basis property” may underlie heterogeneous tuning observed in neural circuits [25], but in our case it imposes modeling limitations: while the random-mode model has configurable spectral properties, it precludes configuration of spatial or anatomical properties.
@@
Because we assume zero-mean mode components (Eq. 4), $\left\langle J_{ij}\right\rangle=0$ for $i\neq j$. Thus, the low-rank structure is not inherited from the expected weight matrix. This differs from the mechanism described by Thibeault et al. [19], who analyze cases where low-rank structure arises because $\left\langle\bm{J}\right\rangle$ is low-rank, with the random part treated as noise. By Weyl’s inequality, the singular values of $\bm{J}$ differ from those of $\left\langle\bm{J}\right\rangle$ by at most the spectral norm of the noise. This mechanism is fundamentally different from the random-mode model’s zero-mean, randomly embedded structure. Which of these mechanisms is more prevalent in real-world networks is an open question.
Alternative generative models. Tiberi et al. [79] proposed a generative model for coupling matrices that parameterizes the eigendecomposition of $\bm{J}$ rather than its SVD. Their formulation does not capture relationships between eigenvalues and eigenvectors. Our SVD-like approach allows simultaneous control over component strengths and their corresponding mode overlaps. Whereas Tiberi et al. [79] studied noise-driven linear networks within a random matrix theory framework, our DMFT formulation applies to nonlinear dynamics.
Connectivity parameterizations resembling the random-mode model have been used to study the correspondence between spiking and rate-based recurrent neural networks [80] and to expedite training spiking networks [81].
Conclusion. The random-mode model and associated analytical techniques provide tools for exploring the relationship between connectivity and collective dynamics. Our parameterization of connectivity is a middle ground between intensive-rank models and i.i.d. random networks. Building on this framework stands to further bridge neuronal connectivity, activity, and function.
Acknowledgements
We thank L. F. Abbott for feedback. A.v.M. was supported by the Swartz Foundation; D.G.C, O.M., and A.L.-K. by the Kavli Foundation and Gatsby Charitable Foundation (grant no. GAT3708); and D.G.C. by the National Institutes of Health (grant no. T32NS064929).

### v3

@@ VII Mode overlaps
Thus far, we have assumed i.i.d. modes, with the factorization $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$ implying that left-right overlaps $\bm{\ell}_{a}^{T}\bm{r}_{a}$ are random and $\mathcal{O}(1/\sqrt{N})$. The same holds for $\bm{\ell}_{a}^{T}\bm{r}_{b}$ with $a\neq b$ due to the more fundamental assumption that $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ factorizes across $a$. However, computations in biological circuits require structured interactions between modes, suggesting stronger, nonrandom overlaps between left and right modes. To test this, we return to the Drosophila connectome and compute the overlap matrix between left and right singular vectors, $O_{ab}=\bm{u}_{a}^{T}\bm{v}_{b}$ [Fig. 6(a)]. Contrary to the small, unstructured overlaps expected under the i.i.d. assumption, the connectome exhibits large, structured overlaps, particularly along the diagonal [Fig. 6(b)].
@@
Figure 6: Right-left mode overlaps and eigenvalue spectra of the fly-brain connectome. (a) Overlap matrix between right and left singular vectors of the connectome. First 500 modes are shown. (b) First 3000 diagonal elements of the overlap matrix from (a). Horizontal lines indicate an approximate cutoff for $\mathcal{O}(1/\sqrt{N})$ random overlaps. (c) Eigenvalue spectra of various connectome-based matrices. Top: actual hemibrain spectrum, exhibiting large outliers along the real axis. Middle: eigenvalues from a realization of the random-mode model (RMM), using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum. Bottom: random-mode model realization with right-left mode overlaps incorporated, where overlaps are given by the inner products of left and right singular vectors from the fly connectome SVD. Data from Scheffer et al. [17], $N=18028$ neurons.
@@
Motivated by the presence of these large diagonal overlaps, we extend the random-mode model to include correlations between corresponding left and right modes. Specifically, we take $P(\ell_{a},r_{a})$ to be a zero-mean bivariate Gaussian with marginal variance $1/N$ and covariance $\rho_{a}/N$, where $\rho_{a}$ is the correlation between the $a$th mode pair. Both $D_{a}$ and $\rho_{a}$ are treated deterministically, and we assume that $M^{-1}\sum_{a}f(D_{a},\rho_{a})$ converges to a limiting value $\left\langle f(D,\rho)\right\rangle_{D,\rho}$ as $N\to\infty$. In this setting, $\bm{\ell}_{a}^{T}\bm{r}_{a}=\rho_{a}$, with negligible $\mathcal{O}(1/\sqrt{N})$ fluctuations around this structured overlap, while $\bm{\ell}_{a}^{T}\bm{r}_{b}$ for $a\neq b$ remains unstructured and $\mathcal{O}(1/\sqrt{N})$. This additional structure produces eigenvalue spectra resembling that of the connectome, including large real outliers [Fig. 6(c)] [68].
@@
This purely diagonal overlap structure, however, can generate only real eigenvalue outliers. The presence of complex outliers in the connectome, together with the full overlap matrix itself [Fig. 6(a)], implies the presence of richer correlations. Capturing such features would require relaxing the full factorization of $P(\{\ell_{a},r_{a}\})$ across $a$ while retaining $\mathcal{O}(M)$ parameter scaling—for example, via blockwise factorization or Markovian correlations across $a$. We leave such generalizations for future work. We also note that by setting $\rho_{a}=1$ for all $a$, the model becomes a Hopfield network storing an extensive set of Gaussian patterns. In such networks, each mode can be “condensed” $\mathcal{O}(1)$ overlap with activity] or “uncondensed” $\mathcal{O}(1/\sqrt{N})$ overlap with activity]. Our analysis assumes no condensed modes, but sufficiently strong positive overlaps break this assumption. The Gaussian case, however, is unable to generate multiple multistable attractors as in a memory system [44]. Note that the form of the single-element density is relevant for the condensed patterns, but not for the uncondensed patterns. Allowing for condensed patterns within our analysis would require specifying this density. See Ref. [69] for a DMFT analysis of generalized Hopfield dynamics in which these ideas are implemented.
@@
In Appendix K, we compute the two- and four-point functions. As in previous cases, the dynamics reduce to single-site processes,
@@
To better understand the self-coupling kernel $R_{\star}(\omega)$, we expand in powers of $\rho$. The coefficients encode connectivity motifs of increasing orders,
@@
We validate this theory using $\alpha=1$ and $D_{a}\propto\exp(-\beta_{D}a/N)$, as before. For overlaps, we use a sigmoidal spectrum for $\rho_{a}$ qualitatively inspired by Fig. 6(b), with multiplicative prefactor $\gamma_{\rho}$ and other parameters chosen such that $\left\langle D\rho\right\rangle_{D,\rho}=0$ (Appendix K.3). Across values of $g_{\text{eff}}$ and $\text{PR}^{D}$, sufficiently strong overlaps, of either sign, reduce the dimension of activity (Fig. 7). The maximum dimension occurs for $\gamma_{\rho}<0$, where dynamics most strongly suppress activity aligned with dominant connectivity modes. This effect weakens as $g_{\text{eff}}$ increases.
In summary, our analysis reveals how connectivity structure shapes collective activity and its dimensionality. When mode overlaps are unstructured, connectivity structure reduces to two parameters: the effective rank $\alpha\text{PR}^{D}$ and effective coupling strength $g_{\text{eff}}$. Structured left-right mode overlaps break this simplification: Collective dynamics then depend on the full joint distribution of component strengths and overlaps, providing an additional mechanism for controlling activity dimensionality.
Figure 7: Effect of right-left mode correlations on dimension of activity in the random-mode model. Dimension of activity $\text{PR}^{\phi}$ versus correlation parameter $\gamma_{\rho}$ in the $(D,\rho)$ joint distribution, for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$. Thin dots, individual simulations; thick dots, means over ten simulations; lines, theoretical predictions. All simulations use $N=5000$ neurons.
@@ VIII.1 Motifs versus global structure
Neuronal connectivity is often described using statistics of motifs: local patterns like chains or cycles among small groups of neurons [73, 22, 74, 75]. These can be estimated from partial observations of a weight matrix which, until recently, set the limits of experimental measurement. Advances in whole-brain connectome reconstruction now allow access to global connectivity, including spectral features. Our parametrization of coupling matrices represents this global structure, with control over both the spectrum and mode overlaps. These spectral features determine how activity modes interact dynamically and thus may more directly dictate neural computations compared to motifs.
@@
That said, in some studies the primary goal is to capture motifs or graph-theoretic features (e.g., degree distributions) over which the random-mode model does not provide control (although some of these features have spectral signatures that could be incorporated). When such features are the focus, other connectivity models are better suited—for instance, maximum-entropy random graph models constrained by fixed structural features (e.g., the configuration model for matching degree distributions).
VIII.1.1 Linking connectivity and activity datasets.
@@
The random-mode model, together with its dynamical theory, could aid in the interpretation of large-scale neural activity recordings in the context of anatomical connectivity, such as in Drosophila, where both the connectome and whole-brain activity data are available [76]. In the current work, we used this connectome to characterize the spectral structure of an actual large neural circuit, stopping short of considering the activity that it generates. Future work could progressively incorporate richer connectome constraints into the model to identify which features most strongly shape collective activity. One possible sequence would be to add the spectrum of component strengths, then diagonal mode overlaps, and finally off diagonal overlaps, assessing at each step how much the correspondence between modeled and observed activity improves. This is similar to approaches in statistical physics that aim to recover higher-order correlations by constraining lower-order statistics [77, 78, 38].
VIII.2 Learning
@@
Learning alters collective dynamics by modifying connectivity, the core principle of machine learning. Our analysis focuses on structured but untrained networks that produce chaotic activity. Nevertheless, the extensive low-rank connectivity structures we study provide a starting point for understanding trained networks.
@@
Large networks trained on simple tasks often show intensive low-rank modifications to their initial couplings, resulting in a finite number of task-relevant eigenvalues [27, 43]. Existing theoretical tools can already describe such systems by taking the trained weights, performing a low-rank decomposition, modeling mode overlaps statistically, and studying the resulting low-rank DMFT [45]. By contrast, empirical studies of large networks trained on real-world tasks suggest the emergence of extensive low-rank structure, with smooth spectral profiles with many significant components [28]. This structure resembles that modeled by the random-mode model. Whether mode overlaps extracted from such networks can be incorporated into the random-mode model framework, and thereby yield a working theory for the extensive-rank case, is an open question.
@@
A more ambitious direction is to develop a DMFT for networks whose weights are learned from data, rather than first training networks and then retrofitting a statistical model. Such a theory would describe not only how data shape the weights, as addressed by classical works [79], but also how data interact with recurrent network dynamics. The path-integral formalism is well suited to this, as it allows derivation of DMFT equations for coupling matrices drawn from a Gibbs ensemble with an energy function given by a task loss. While this approach has been applied to low dimensional tasks [80], applying it to high-dimensional tasks could yield a theory of how such tasks generate extensive-rank weights that, in turn, produce extensive-dimensional activity whose single-neuron and collective properties the theory would describe.
@@
Another route to extensive-rank structure is via networks that perform many tasks, which could be modeled, for example, by representing mode overlaps through blockwise factorization of $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ across $a$. This multitask scenario is increasingly relevant in both neuroscience and machine learning [81, 82].
@@
Even in standard recurrent neural network training, initializing couplings with low-rank structure could improve procedures such as FORCE learning, which rely on suppressing chaos [12, 13]. Lower-dimensional attractors may be easier to control, potentially leading to faster convergence.
VIII.3 Modeling functional consequences of single-neuron heterogeneity
@@
A central question in neuroscience is how neuronal diversity shapes collective dynamics and computation [83]. Appendix J presents a general framework that assigns each neuron a parameter vector $\bm{\theta}$ and enables the study of how different forms of heterogeneity affect properties such as dimensionality. Section VI focuses on gain heterogeneity to model firing-rate diversity, but this framework could also incorporate more intricate cell-specific features, such as electrophysiological characteristics [84], linking single-neuron heterogeneity to functional consequences for population activity.
VIII.4 Transient dynamics and excitation-inhibition
@@
Our analysis has focused on stationary states with translation-invariant temporal correlations. Many neural computations, however, rely on transient dynamics that break stationarity and depend on absolute time. Extending the DMFT accordingly is numerically tractable for two-point functions but would become unwieldy for four-point functions, $\Psi^{a}_{\star}(t_{1},t_{2},t_{3},t_{4})$. Representing and manipulating such a large object may require approximations or low-rank decomposition.
@@
One notable transient phenomenon is transient amplification, linked to non-normal coupling matrices that naturally arise from excitatory-inhibitory segregation (Dale’s law) [85, 86]. Although the random-mode model does not enforce Dale’s law, the resulting matrices are nevertheless non-normal and may exhibit transient amplification; such effects may be amplified through structured off diagonal overlaps between left and right modes. Non-normality also decouples the highest-variance dimensions from the slowest modes. Relatedly, extending the random-mode model to respect Dale’s law and incorporating a nonsaturating $\phi(x)$—achieving stability via excitation-inhibition balance rather than saturation [87, 88]—could enable more direct comparisons to both connectivity and activity data.
VIII.5 Random basis property
@@
The random-mode model assumes that the $2M$ mode components associated with each neuron are sampled i.i.d. for each neuron [Eq. (3)]. This leads to neuronal permutation symmetry of the distribution over $\bm{J}$. Furthermore, when the marginal distributions over left and right mode components are isotropic Gaussians, as we have assumed, the embedding into neuronal space is, by definition, random and Gaussian. This “random basis property” may underlie heterogeneous tuning observed in neural circuits [29], but in our case it imposes modeling limitations: While the random-mode model has configurable spectral properties, it precludes configuration of spatial or anatomical properties.
@@
Because we assume zero-mean mode components [Eq. (4)], $\left\langle J_{ij}\right\rangle=0$ for $i\neq j$. Thus, the low-rank structure is not inherited from the expected weight matrix. This differs from the mechanism described by Thibeault et al. [23], who analyze cases where low-rank structure arises because $\left\langle\bm{J}\right\rangle$ is low rank, with the random part treated as noise. By Weyl’s inequality, the singular values of $\bm{J}$ differ from those of $\left\langle\bm{J}\right\rangle$ by, at most, the spectral norm of the noise. This mechanism is fundamentally different from the random-mode model’s zero-mean, randomly embedded structure. Which of these mechanisms is more prevalent in real-world networks is an open question.
VIII.6 Alternative generative models
Tiberi et al. [89] proposed a generative model for coupling matrices that parametrizes the eigendecomposition of $\bm{J}$ rather than its SVD. Their formulation does not capture relationships between eigenvalues and eigenvectors. Our SVD-like approach allows simultaneous control over component strengths and their corresponding mode overlaps. Whereas Tiberi et al. [89] studied noise-driven linear networks within a random-matrix-theory framework, our DMFT formulation applies to nonlinear dynamics.
Connectivity parametrizations resembling the random-mode model have been used to study the correspondence between spiking and rate-based recurrent neural networks [90] and to expedite training spiking networks [91].
IX Conclusion
The random-mode model and associated analytical techniques provide tools for exploring the relationship between connectivity and collective dynamics. Our parametrization of connectivity is a middle ground between intensive-rank models and i.i.d. random networks. Building on this framework stands to further bridge neuronal connectivity, activity, and function.
Acknowledgments
We thank L. F. Abbott for feedback. A. v. M. was supported by the Swartz Foundation and the Kempner Institute for the Study of Natural and Artificial Intelligence. D. G. C, O. M., and A. L.-K. were supported by the Kavli Foundation and Gatsby Charitable Foundation (Grant No. GAT3708). D. G. C. was also supported by the National Institutes of Health (Grant No. T32NS064929). A. L.-K. was also supported by the National Science Foundation (Grant No. 2443158) and the National Institutes of Health (Grant No. RF1DA060772).